The average life of a bread-making machine is 7 years, with a standard deviation of 1 year. Assuming that the lives of these machines follow approximately a normal distribution, find the value of the sample mean to the right of which 15% of the means computed from random samples of size 9 would fall.


1
Expert's answer
2022-01-10T18:37:34-0500

"P(Z>z)=0.15."

"P(Z<z)=0.85."

"z=1.04."

"\frac{\bar x-7}{\frac{1}{\sqrt{9}}}=1.04."

"\bar x=7+\frac{1.04}{3}=7.35."